OpenStudy (anonymous):

Is there a real solution to -y = x^2 + 9 and y = -1?

OpenStudy (opcode):

$$-y = x^2 + 9 \implies x^2 + 9 = - 1$$ $$\implies x^2 = - 10$$ What can you determine from that?

OpenStudy (anonymous):

Since there is no square root for -10 there is no real solution? @Opcode

OpenStudy (opcode):

You are correct. :-) There is no real solution.

OpenStudy (anonymous):

Thank you

OpenStudy (opcode):

No problem. :-)